5. Majority-rule benefit-cost analysis

Walter, Yvonne, Jared, and Jessica are college roommates. They're trying to decide where the four of them should go for spring break: Miami or Virginia Beach. If they order the tickets by 11:00 PM on February 1, the cost will be just $500 per person. If they miss that deadline, the cost rises to $1,200 per person.

The following table shows the benefit (in dollar terms) that each roommate would get from the two trips.

Roommate Benefit from Miami
Walter $950 $450
Yvonne $550 $750
Jared $650 $650
Jessica $750 $950

The roommates tend to put off making decisions. So, when February 1 rolls around and they still haven't made a decision, they schedule a vote for 10:00 PM that night. In case of a tie, they will flip a coin between the two vacation destinations.

The roommates will get the most total benefit if they choose to go to ______.

Given the individual benefits each roommate receives from the two trips, which trip will each roommate vote for?

Fill in the table with each roommate's preferred location, assuming that a given roommate will abstain if he or she has no preference.

Roommate Vote
Walter _____
Yvonne _____
Jared _____
Jessica _____

Under majority rule, the roommates will vote to go to __________  . Therefore, majority rule leads to an economically ________ outcome.

Suppose Jessica misses the vote, leaving Walter, Yvonne, and Jared to figure out where they're going to go.

Walter and Yvonne argue for their preferred destinations, but Jared offers to vote with Walter if Walter will vote on his side in an upcoming class election. This is an example of ________
Answer #1
Roomate Net Benefit from Miami Net benefit from Virginia
Walter (950-500)=$450 (450-500)=$(50)
Yvonne (550-500)=$50 (750-500)=$250
Jared (650-500)=$150 (650-500)=$150
Jessica (750-500)=$250 (950-500)=$450
Total 900 800
Net benefit per person 900/4= $225 per person 800/4 = $200 per person

1) The roommates will get total benefit if they choose to go to Miami .

2)Table of voting :-

Roommate Vote
Walter Miami
Yvonne Virginia island
Jared Any ( if considers net benefit then Miami )
Jessica Virginia island


3) Under Majority rule roommates will vote to go to Virginia islands , thus the Majority rule leads to an economically unviable outcome

4) Jared offers to vote for Walter if Walter votes for him in the class election . This is an example of Biased voting due to comparative advantage.